OpenStudy (anonymous):

y varies jointly with p and q. When p = 10 and q = 2, y = 60. Find y when p = 5 and q = 5 y = 300 y = 250 y = 100 y = 75

jhonyy9 (jhonyy9):

75

OpenStudy (alfie):

I'd say, y = 75.... I am not completely sure though. Nevertheless...!

OpenStudy (alfie):

y= 3pq

OpenStudy (anonymous):

yeah according to me also 75

OpenStudy (anonymous):

(10*2)*3=60 similarily, (5*5)*3=75

OpenStudy (anonymous):

\(y\) varies jointly with \(p\) and \(q\) mean \(y=kpq\), where k is some constant needed to be found. We have \(60=k(10)(2) \implies k=3\). That means that our equation would be \(y=3pq\). So at \(p=q=5\), \(y=3(5)(5)=75\).